# IndicSentEval: How Effectively do Multilingual Transformer Models encode Linguistic Properties for Indic Languages?

### v1

@@ 7 Experimental Results
We measure probing accuracy independently for each multilingual model, within each layer separately. Along with six Indic languages, we measure the probing accuracy for English language across all multilingual models and compare the findings of Indic languages against English. Further, we investigate the robustness of these multilingual models by perturbing the input sentences.
@@ 7.1 Probing Results
How effectively do multilingual models encode hierarchy of linguistic structure for Indic languages? We assess the linguistic structure by contrasting universal models trained on 100+ languages with those specifically trained on Indic languages. Unless otherwise specified, the results presented in the main paper reflect an average accuracy across encoder-based universal models, decoder-based universal models and Indic-specific models.
@@
Surface-level tasks. We show the accuracy scores obtained for surface level task (i.e. SentLen) in the first row in Fig. 2. Analyzing the performance across languages, we observe the following patterns: (i) For encoder-based universal models as well as for Indic models, there is a trend of higher accuracy in the early (or lower) layers, which decreases in the later (or higher) layers. This pattern is expected, as surface-level tasks generally require minimal processing. (ii) Notably, the decoder-based universal multilingual models deviates from this trend, showing lower accuracy in the early layers and higher accuracy in the later layers. This unusual pattern in decoder-based universal models could be attributed to their unique autoregressive architectural design, and their approach to representing languages, especially those less common in the dataset. This suggests that decoder-based universal models may require their deeper layers to effectively grasp the nuances of these languages, even for tasks at the surface level. (iii) Overall, among all the models, Indic models report best accuracy, while encoder-based universal models display poorer performance. This is likely because Indic models are specifically trained on Indic languages, making them more attuned to the nuances and idiosyncrasies of these languages. On the other hand, universal models, which are designed for universal applicability, might struggle with specific linguistic features unique to Indic languages. These features include script differences, morphological complexity, and visual factors such as orthography and word length.
@@
We present the individual model specific results across language in Appendix Figs. 5 to 10. We observe that among all the models, IndicBERT reports best accuracy, while universal multilingual model XLM-R displays poor performance. More detailed analysis is reported in Appendix D.
@@
Figure 3: Probing task results: Layerwise accuracy comparisons between various multilingual representations on semantic probing tasks. For Malayalam, there is an absence of SSF data for the VerbGen, VerbPer, and VerbNum tasks. We report the layerwise probing accuracies for individual multilingual models in Appendix Figs. 5 to 10.
@@
Syntactic tasks. In the bottom two rows of Fig. 2, we display the accuracy scores for syntactic tasks. For TreeDepth, we observe that probing accuracy tends to be higher in the middle layers for various (model, language) combinations. This trend is particularly notable in both encoder-based universal models and Indic models. Moreover, this pattern is consistent for three languages: hi, kn and ml. However, decoder-based multilingual models do not exhibit any clear layer-wise trend, and the same applies to the other three Indic languages. This suggests a model-specific and language-specific affinity in handling syntactic complexity, where certain models are more adept at processing syntactic information in specific layers, and this proficiency varies across different languages. In contrast to TreeDepth, for BShift task, we generally observe higher probing accuracy in the later layers for both encoder-based universal and Indic models across various languages. Notably, decoder-based universal models exhibit a decreasing trend in accuracy for kn, ml and te, while showing an increasing trend for hi and ur. This suggests that different models’ layers may specialize in different types of syntactic processing, with some models better handling tasks like BShift in their later layers. Overall, when comparing performance across models and tasks, decoder-based universal models stand out for their superior accuracy in capturing tree depth information across different languages. Conversely, Indic models show notable proficiency in BShift. These distinctions highlight how different models may be better suited for different types of syntactic analysis. These observations contribute to a deeper understanding of how various multilingual models process syntactic information, demonstrating both model-specific and language-specific trends and capabilities in linguistic tasks.
@@
From the individual model specific results in Appendix Figs. 5 to 10, we observe that among all the models, InfoXLM and mT5 stand out for their superior accuracy in capturing tree depth information across different languages. Conversely, Indic model MuRIL shows notable proficiency in BShift. More detailed analysis is reported in Appendix D.
@@
From SubjNum and ObjNum results, we make the following observations. For decoder-based universal models, we observe an increasing trend from lower to higher layers for hi, kn and te. This suggests that for these languages, the models become more proficient in handling semantic tasks related to SubjNum and ObjNum as we move to the higher layers. When considering other languages and models, we note that both encoder-based and Indic models exhibit an increasing trend for the ObjNum task, specifically for hi and mr. This highlights the variability in model performance based on the language. Interestingly, the middle layers of the encoder-based universal models and Indic models show higher probing accuracy for languages such as hi, ml, mr and te in the SubjNum task. For the verb-related tasks, decoder-based universal, performs the best for most languages and tasks. Also, in most cases, the last layer is the most predictive, except in mr and ur for gender, number and person detection, where initial layers provide better results. This indicates that gender, number and person detection in mr and ur is straightforward and does not need deep processing. We observe this decrease trend for encoder-based universal models for language te. For the Indic models, we observe an increasing trend for two languages hi and kn across tasks. Overall, both encoder-based universal and Indic models show mixed trend in performance across layers.
@@
From the individual model specific results in Appendix Figs. 5 to 10, we observe that XLM-R exhibits lower accuracy and lacks a discernible trend when it comes to capturing semantics. MuRIL has the highest probing accuracy although it performs the worst for kn ObjNum task. More detailed analysis is reported in Appendix D.
@@
Comparison of encoding performance of multilingual models for linguistic properties for English vs Indic languages. For this comparison, we conduct probing for English across nine multilingual models for five tasks: SentLen, TreeDepth, BShift, SubNum, and ObjNum using SentEval dataset Conneau et al. (2018). Fig. 11 in the Appendix reports the probing accuracy. Surprisingly, across all multilingual models, we observe that surface features show a decreasing trend from lower to higher layers. For syntactic TreeDepth feature probing accuracy is higher in the middle layers, while BShift has an increasing trend from lower to higher layers. Finally, the semantic tasks SubNum and ObjNum are best encoded in the later layers. This implies that the English language is encoded in the same manner in both universal and Indic multilingual models, whereas Indic languages show mixed results across models.
@@
Overall insights from probing experiments. While Indic-specific models like MuRIL and IndicBERT are likely the best at capturing language properties within the realm of Indic languages due to their targeted training, both encoder-based and decoder-based universal models like mBERT, InfoXLM, BLOOM and mGPT show mixed results. Their broader training might enable them to capture more general properties across many languages, but they may lack a deep understanding specific to each language, particularly those less represented in their training corpus. The effectiveness of these models thus depends on the specific linguistic features and tasks being evaluated, as well as the range of languages being considered.
Overall insights from probing experiments. While Indic-specific models like MuRIL and IndicBERT are likely the best at capturing language properties within the realm of Indic languages due to their targeted training, both encoder-based and decoder-based universal models like mBERT, InfoXLM, BLOOM and mGPT show mixed results. Their broader training might enable them to capture more general properties across many languages, but they may lack a deep understanding specific to each language, particularly those less represented in their training corpus. The effectiveness of these models thus depends on the specific linguistic features and tasks being evaluated, as well as the range of languages being considered.
We believe that we obtain mixed results for Indic languages for the following reasons. Universal multilingual models with broader training might enable them to capture more general properties across many languages, but they may lack a deep understanding specific to each language, particularly those less represented in their training corpus. The effectiveness of these models thus depends on the specific linguistic features and tasks being evaluated, as well as the range of languages being considered.
Overall insights for Indic languages mBERT and MuRIL capture linguistic features for hi very well. For mr and te, mT5 and MuRIL show better accuracy for surface, syntactic and semantic tasks. This is in alignment with the fact that hi, mr, and te are better represented in pretraining datasets for these models.
7.2 Perturbation Results
We perform 13 different text perturbations to understand the contribution of specific word types and sentence structures to the encoding capabilities of multilingual language models [^3] Perturbations don’t make sense for BShift probing task.. We analyze the impact of such text perturbations for every pair of (model, language), (perturbation, model), (language, probing task), (model, probing task), (perturbation, language) and (perturbation, probing task) in Tables 23456 and 7, respectively. All these tables show weighted averages across marginalized dimensions.
@@ 7.2 Perturbation Results
mBERT Indic XLM-R Info MuRIL BLOOM mT5 mGPT XGLM
@@
Sent Tree Subj Obj Verb Verb Verb
Len Depth Num Num Gen Num Per
mBERT 0.398 0.487 0.916 0.924 0.839 0.903 0.939
IndicBERT 0.364 0.504 0.928 0.931 0.874 0.947 0.969
XLM-R 0.492 0.564 1.030 0.996 0.968 1.001 1.006
InfoXLM 0.855 0.836 0.895 0.916 0.738 0.808 0.926
MuRIL 0.385 0.471 0.900 0.935 0.833 0.906 0.918
BLOOM 0.604 0.870 0.933 0.930 0.892 0.849 0.899
mT5 0.582 0.905 0.915 0.931 0.898 0.853 0.914
mGPT 0.918 0.872 0.932 0.936 0.820 0.879 0.931
XGLM 0.761 0.865 0.930 0.925 0.834 0864 0.904
@@
Which multilingual models are more robust across probing tasks? Table 5 shows that universal models demonstrate greater robustness compared to Indic models and mBERT. Additionally, surface and syntactic probing tasks are significantly impacted by perturbations compared to semantic properties.
Which text perturbations have the greatest impact on probing tasks across six Indic languages? Tables 6 and 7 show that dropping nouns and verbs significantly affects accuracy across all six languages. Specifically, dropping verbs impacts accuracy of three verb-based semantic tasks. Similarly, position perturbations, such as keeping nouns in specific positions, affect verb tasks, while keeping verbs or nouns affects surface and syntactic tasks.
Which layers are more effected due to text perturbations for Indic languages? From Table 8, we observe: (i) across all Indic languages, early layers are impacted for surface properties. (ii) for TreeDepth syntactic property, middle layers are affected for all languages except te. (iii) Surprisingly, later layers are impacted more than early and middle layers. Regarding semantic properties, for ur, early to middle layers are impacted more than later layers, except for ObjNum. Similarly, for kn, SubjNum and ObjNum are impacted on middle layers; for mr, VerbNum and VerbPer are affected.
@@
hi kn ml mr te ur
SentLen 1,2,3 4,3,2 3,4,1 1,3,2 1,2 3,1,2
TreeDepth 7,5,6 5,4,9 4,5,3 4,8,9 10,11,12 3,2,5
VerbGen 10,8,5 11,4,12 - 11,10,12 6,8,9 3,8,9
VerbNum Equal 3,4,5 - 2,4,5 10,12 5,4,3
VerbPer Equal Equal - 3,4,5 5,11,12 4,5
SubjNum N/A Equal 9,8,7 7,10,9 10,11 1,2
ObjNum 11,12 4,3,9 Equal 11,10,12 11,12 10,9,11
@@ 8 Discussion and Conclusion
We investigated how well multilingual Transformer-based models comprehend linguistic structures in 6 Indic languages by evaluating 9 models using 8 probing tasks. To support this study, we contributed the IndicSentEval dataset. Our results show that Indic-specific models like MuRIL and IndicBERT are the most effective at capturing linguistic properties within Indic languages due to their targeted training. Universal models like mBERT, InfoXLM, BLOOM, mGPT and XGLM exhibit mixed results. Our perturbation analysis reveals that decoder-based models are the most robust, as also observed by Neerudu et al. (2023). Verbs and the order of words are the most important signals that help the models encode linguistic structures. The TreeDepth is the most sensitive to all mentioned text perturbations, while SubjNum and ObjNum are the most resilient.
Overall, our study represents the first analysis of the interpretability of both Universal and Indic multilingual language models across six Indic languages where several languages have large training corpora while some have less. Our scientific findings from this model interpretability analysis via both probing and perturbations shed light on how language models capture language hierarchy and how training data influences the language understanding across layers in these models. In the probing analysis, Indic-specific models like MuRIL and IndicBERT are likely the best at capturing language hierarchy while Universal models like mBERT, InfoXLM, BLOOM, mGPT and XGLM show mixed results. Surprisingly, universal models show greater resilience to perturbations in at least four Indic languages. In contrast, the universal model (mBERT) and the Indic-specific models (IndicBERT and MuRIL) display a more significant accuracy drop across all the Indic languages. This suggests the necessity for multilingual models that are robust to perturbations and can capture language hierarchy regardless of their training data. Overall, our findings demonstrate significant variability in the ability of current multilingual models to capture surface, syntactic, and semantic structures across different Indic languages. A hierarchy of language proficiency is discernible primarily for languages with more extensive training datasets. This highlights the necessity for innovative approaches in multilingual modeling that can accurately capture language structures, even in low-resource languages.
Our findings have direct implications for the emerging field of language model interpretability. Language proficiency is observed only for models with more comprehensive training datasets. The encoding ability and impact of perturbations varies across different languages and models, indicating the need for robust training strategies and architectures tailored to handle linguistic variations and perturbations effectively.

### v2

@@ 5 Experimental Results
We measure probing accuracy independently for each multilingual model, within each layer separately. Along with six Indic languages, we measure the probing accuracy for English language across all multilingual models and compare the findings for Indic languages against English. Further, we investigate the robustness of these multilingual models by perturbing the input sentences.
@@ 5.1 Probing Results
How effectively do multilingual models encode hierarchy of linguistic structure for Indic languages? We assess the linguistic structure by contrasting universal models trained on 100+ languages with those specifically trained on Indic languages only. Unless otherwise specified, the results presented in the main paper reflect an average accuracy across encoder-based universal models, decoder-based universal models and Indic-specific models.
@@
Surface-level tasks. We show the accuracy scores obtained for surface level task (i.e. SentLen) in the first row in Fig. 2. Analyzing the performance across languages, we observe the following patterns: (i) For encoder-based universal models as well as for Indic models, there is a trend of higher accuracy in the early (or lower) layers, which decreases in the later (or higher) layers. This pattern is expected, as surface-level tasks generally require minimal processing. (ii) Notably, the decoder-based universal multilingual models deviates from this trend, showing lower accuracy in the early layers and higher accuracy in the later layers. This seemingly unusual pattern in decoder-based universal models is actually intuitive because masked self attention in their autoregressive architectural design implies that only deeper layers in such models can effectively grasp the input length. (iii) Overall, among all the models, Indic models report best accuracy, while encoder-based universal models display poorer performance. This is likely because Indic models are specifically trained on Indic languages, making them more attuned to the nuances and idiosyncrasies of these languages. On the other hand, universal models, which are designed for universal applicability, might struggle with specific linguistic features unique to Indic languages. These features include script differences, morphological complexity, and visual factors such as orthography and word length. Tokenizers of universal models tokenize Indic language inputs to many tokens with little correlation with actual input length in words.
@@
We present the individual model-specific results across languages in Figs. 5 to 10 in Appendix G. We observe that among all the models, IndicBERT reports the best accuracy, while universal multilingual model XLM-R displays poor performance. More detailed analysis is reported in Appendix G.
@@
Figure 3: Probing task results: Layerwise accuracy comparisons between various multilingual representations on semantic probing tasks. For Malayalam, there is an absence of SSF data for the VerbGen, VerbPer, and VerbNum tasks. We report the layerwise probing accuracies for individual multilingual models in Figs. 5 to 10 in Appendix G.
@@
Syntactic tasks. In the bottom two rows of Fig. 2, we display the accuracy scores for syntactic tasks. For TreeDepth, we observe that probing accuracy tends to be higher in the middle layers for various (model, language) combinations. This trend is particularly notable in both encoder-based universal models and Indic models. Moreover, this pattern is consistent for three languages: hi, kn and ml. However, decoder-based multilingual models do not exhibit any clear layer-wise trend, and the same applies to the other three Indic languages. Encoders, with their bidirectional attention, might be inherently better at capturing hierarchical structures, while decoders, with their unidirectional attention, might excel in tasks requiring sequential processing. The lack of a clear trend in decoder-based multilingual models might indicate that these models distribute syntactic processing more evenly across layers. The varying performance across languages underscores the importance of considering linguistic diversity in model training.
In contrast to TreeDepth, for BShift task, we generally observe higher probing accuracy in the later layers for both encoder-based universal and Indic models across various languages. Notably, decoder-based universal models exhibit a decreasing trend in accuracy for kn, ml and te, while showing an increasing trend for hi and ur. This suggests that languages with different syntactic structures may require different layers to process syntactic tasks effectively.
Overall, when comparing performance across models and tasks, decoder-based universal models stand out for their superior accuracy in capturing tree depth information across different languages. Conversely, Indic models show notable proficiency in BShift. These findings highlight the need for comprehensive evaluation across multiple tasks and languages to understand model capabilities fully. Evaluating models on diverse syntactic tasks can reveal strengths and weaknesses that might not be apparent from a single task. These observations contribute to a deeper understanding of how various multilingual models process syntactic information, demonstrating both model-specific and language-specific trends and capabilities in linguistic tasks.
@@
From the individual model specific results in Figs. 5 to 10 in Appendix G, we observe that among all the models, InfoXLM and mT5 stand out for their superior accuracy in capturing tree depth information across different languages. Conversely, Indic model MuRIL shows notable proficiency in BShift. More detailed analysis is reported in Appendix G.
@@
From SubjNum and ObjNum results, we make the following observations. For decoder-based universal models, we observe an increasing trend from lower to higher layers for hi, kn and te. Dravidian languages often have more complex morphological systems for marking plurality, with a variety of suffixes and sometimes even changes in the noun stem itself. Hence, this increasing trend makes sense for kn and te. When considering other languages and models, we note that both encoder-based and Indic models exhibit an increasing trend for the ObjNum task, specifically for hi and mr. This highlights the variability in model performance based on the language. Interestingly, the middle layers of the encoder-based universal models and Indic models show higher probing accuracy for languages such as hi, ml, mr and te in the SubjNum task.
For the verb-related tasks, decoder-based universal models perform the best for most languages and tasks. Also, in most cases, the last layer is the most predictive, except in mr and ur for gender, number and person detection, where initial layers provide better results. This indicates that gender, number and person detection in mr and ur is straightforward and does not need deep processing. We observe a decreasing trend for encoder-based universal models for language te. For the Indic models, we observe an increasing trend for two languages hi and kn across tasks. Overall, both encoder-based universal and Indic models show mixed trend in performance across layers.
@@
From the individual model specific results in Figs. 5 to 10 in Appendix G, we observe that XLM-R exhibits lower accuracy and lacks a discernible trend when it comes to capturing semantics. MuRIL has the highest probing accuracy although it performs the worst for kn ObjNum task. More detailed analysis is reported in Appendix G.
@@
Comparison of encoding performance of multilingual models for linguistic properties for English vs Indic languages. We conduct probing for English across nine multilingual models for five tasks: SentLen, TreeDepth, BShift, SubjNum, and ObjNum using SentEval dataset (Conneau et al., 2018). Fig. 11 in Appendix G reports the probing accuracy. Surprisingly, across all multilingual models, we observe that surface features show a decreasing trend from lower to higher layers. For syntactic TreeDepth feature probing accuracy is higher in the middle layers, while BShift has an increasing trend from lower to higher layers. Finally, the semantic tasks SujbNum and ObjNum are best encoded in the later layers. This implies that English is encoded in the same manner in both universal and Indic multilingual models, whereas Indic languages show mixed results across models.
@@
Overall insights from probing experiments. While Indic-specific models like MuRIL and IndicBERT are likely the best at capturing language properties within the realm of Indic languages due to their targeted training, both encoder and decoder-based universal models like mBERT, InfoXLM, BLOOM and mGPT show mixed results. Their broader training might enable them to capture more general properties across many languages, but they may lack a deep understanding specific to each language, particularly those less represented in their training corpus. The effectiveness of these models thus depends on specific linguistic features and tasks, as well as the range of languages being considered. mBERT and MuRIL capture linguistic features for hi very well. For mr and te, mT5 and MuRIL show better accuracy for surface, syntactic and semantic tasks. This is in line with the fact that hi, mr, and te are better represented in pretraining datasets for these models.
@@
Sent Tree Subj Obj Verb Verb Verb
Len Depth Num Num Gen Num Per
mBERT 0.398 0.487 0.916 0.924 0.839 0.903 0.939
IndicBERT 0.364 0.504 0.928 0.931 0.874 0.947 0.969
XLM-R 0.492 0.564 1.030 0.996 0.968 1.001 1.006
InfoXLM 0.855 0.836 0.895 0.916 0.738 0.808 0.926
MuRIL 0.385 0.471 0.900 0.935 0.833 0.906 0.918
BLOOM 0.604 0.870 0.933 0.930 0.892 0.849 0.899
mT5 0.582 0.905 0.915 0.931 0.898 0.853 0.914
mGPT 0.918 0.872 0.932 0.936 0.820 0.879 0.931
XGLM 0.761 0.865 0.930 0.925 0.834 0.864 0.904
5.2 Perturbation Results
We perform 13 different text perturbations to understand the contribution of specific word types and sentence structures to the encoding capabilities of multilingual language models [^2] Perturbations do not make sense for BShift probing task.. We analyze the impact of such text perturbations for every pair of (model, language), (model, probing task), (perturbation, model) and (language, probing task) in Tables 334 and 5 respectively.
We report results for the layerwise perturbation analysis and other pairs like (perturbation, language) and (perturbation, probing task) in Tables 67 and 8. All these tables show weighted averages across marginalized dimensions.
Which multilingual models are more robust across probing tasks? Table 3 shows that universal models have greater robustness compared to Indic models and mBERT. Additionally, surface and syntactic probing tasks are significantly impacted by perturbations compared to semantic properties.
mBERT
Indic
XLM-R
Info
MuRIL
BLOOM
mT5
mGPT
XGLM
@@ 5.2 Perturbation Results
Which layers are more affected due to text perturbations for Indic languages? Table 6, we observe: (i) across all Indic languages, early layers are impacted for surface properties. (ii) for TreeDepth syntactic property, middle layers are affected for all languages except te. (iii) Surprisingly, later layers are impacted more than early and middle layers. Regarding semantic properties, for ur, early to middle layers are impacted more than later layers, except for ObjNum. Similarly, for kn, SubjNum and ObjNum are impacted on middle layers; for mr, VerbNum and VerbPer are affected.
Which text perturbations have the greatest impact on probing tasks across six Indic languages? Tables 7 and 8 show that dropping nouns and verbs significantly affects accuracy across all six languages. Specifically, dropping verbs impacts accuracy of three verb-based semantic tasks. Similarly, position perturbations, such as keeping nouns in specific positions, affect verb tasks, while keeping verbs or nouns affects surface and syntactic tasks.
hi kn ml mr te ur
SentLen 1, 2, 3 4, 3, 2 3, 4, 1 1, 3, 2 1, 2 3, 1, 2
TreeDepth 7, 5, 6 5, 4, 9 4, 5, 3 4, 8, 9 10, 11, 12 3, 2, 5
VerbGen 10, 8, 5 11, 4, 12 - 11, 10, 12 6, 8, 9 3, 8, 9
VerbNum Equal 3, 4, 5 - 2, 4, 5 10, 12 5, 4, 3
VerbPer Equal Equal - 3, 4, 5 5, 11, 12 4, 5
SubjNum N/A Equal 9, 8, 7 7, 10, 9 10, 11 1, 2
ObjNum 11, 12 4, 3, 9 Equal 11, 10, 12 11, 12 10, 9, 11
@@
5.3 Correlation Analysis of Probing with Downstream Tasks
We conducted a correlation analysis between our IndicSentEval probing results and downstream task performance from the IndicGLUE benchmark. Specifically, we analyze how the performance of the IndicSentEval probing tasks correlates with the results of the IndicGLUE downstream tasks and other benchmarks. This includes models such as IndicBERT, MuRIL, mBERT, and XLM-R across syntactic and semantic tasks like NER, POS tagging, sentiment analysis, and natural language inference. The Table 9 shows the aggregated scores across multiple Indic languages. We make the following observations from Table  9: (i) Models like MuRIL and IndicBERT, which scored highest on syntactic and semantic probes, also performed best on corresponding tasks like POS tagging, NER (syntactic), and sentiment classification, NLI (semantic). (ii) Universal models like mBERT showed weaker probe performance and correspondingly lower scores on most IndicGLUE tasks, especially for morphologically rich languages. (iii) Tasks requiring deeper understanding (e.g. NLI) correlated more strongly with semantic probes, while syntactic probes aligned better with tagging tasks. This implies alignment between what probes measure and what different task categories demand. These findings suggest that the linguistic properties measured by probing are indeed predictive of real-world task performance, particularly in morphologically rich Indic languages.
Model Syntactic Semantic POS NER Sentiment NLI
Probes Probes Tagging Analysis
IndicBERT 0.82 0.80 0.88 0.84 0.81 0.79
MuRIL 0.85 0.86 0.89 0.83 0.85 0.87
mBERT 0.72 0.70 0.76 0.88 0.70 0.69
XLM-R 0.80 0.78 0.85 0.85 0.83 0.82
Table 9: Performance comparison of multilingual models across various probing and downstream tasks.
@@ 6 Discussion and Conclusion
We evaluated 9 multilingual Transformer-based models on 8 probing tasks to understand linguistic structures in 6 Indic languages, using our contributed IndicSentEval dataset. Indic-specific models like MuRIL and IndicBERT excel due to targeted training, while universal models like mBERT, InfoXLM, BLOOM, mGPT, and XGLM show mixed results. Perturbation analysis reveals decoder-based models are the most robust. Overall, our findings highlight the importance of language model interpretability. Language proficiency is seen in models with comprehensive training datasets. Encoding ability and perturbation impact vary across languages and models, underscoring the need for robust training strategies and tailored architectures to handle linguistic variations and perturbations effectively.